OpenStudy (anonymous):

prove: (sinx/1-sinx)-(sinx/1+sinx)=2tan^2x

OpenStudy (anonymous):

First thing when proving an identity is to only work on one side. The easiest side to work on (in my opinion is the left side. Rewrite the left side using the LCD, and combine like terms. You should now have a new fraction with the denominator being a trig identity. You can subsititute into the denominator and then your will have another identity (the whole fraction) which will be equal to the right side of the identity.

OpenStudy (bibby):

Get like denominators and simplify. Remember that\[\sin^2(x)+\cos^2(x) = 1\]and\[\frac{sinx}{cosx}=tanx\]